# Reliability, and Global Privacy

### v3

@@ Abstract
This paper studies privacy-sensitive federated learning (FL) under unreliable communication, with a focus on secure aggregation and straggler mitigation. To preserve user privacy without compromising the utility of the global model, secure aggregation emerges as a promising approach by coordinating the use of privacy-preserving noise (secret keys) across participating clients. However, the unreliable communication will randomly disrupt the key coordination and disable the exact recovery of the global model in secure aggregation. Furthermore, unreliable communication can distort the optimization trajectory, causing the global model to deviate further from the intended global optimum. To address these challenges, we propose Secure Cooperative Gradient Coding (SecCoGC), a practical solution that achieves accurate aggregation with arbitrarily strong privacy guarantees and is inherently robust to communication uncertainties. SecCoGC operates natively in the real field, making it directly deployable in real-world FL systems. To ensure fairness in privacy protection, we further introduce Fair-SecCoGC, an extension of SecCoGC that enforces equitable privacy preservation across all clients. Notably, Fair-SecCoGC achieves optimal privacy under a per-key total power constraint. We formally formulate the problem of secure aggregation in the real field and present both general and computationally efficient methods for secret key construction. Our privacy analysis covers both Local Mutual Information Privacy (LMIP) and Local Differential Privacy (LDP) across all protocol layers, accounting for intermittent networks and correlation among secret keys. In addition, we characterize the system reliability and convergence properties of the proposed scheme. Experimental results demonstrate that SecCoGC achieves strong resilience to unreliable communication while maintaining arbitrarily strong privacy guarantees, yielding test accuracy improvements of $20\%$ to $70\%$ over existing privacy-preserving methods, without requiring any prior information.
Index Terms:
Federated learning, Unreliable communication, Secure Aggregation, Straggler Mitigation, Fairness, Local Mutual Information Privacy, Local Differential Privacy.
@@ I Introduction
I-A Literature Review
Distributed learning (DL) mitigates the computational bottleneck in conventional centralized machine learning (ML) by leveraging edge device (client) capabilities. However, decentralization incurs significant communication overhead during the transmission of datasets. To overcome the communication bottleneck in DL, federated learning (FL) enables clients to collaboratively train models based on their local datasets, thus significantly reducing communication costs and inherently enhancing user privacy by keeping training dataset local [1, 2].
Despite the enhanced privacy provided by restricting raw data sharing, FL systems remain vulnerable to indirect model-based privacy leakage. Even without direct access to the raw datasets, adversaries can still infer sensitive information from the local model updates, making FL susceptible to various model-based attacks, such as inversion attacks (IA) and membership inference attacks (MIA). In IA, adversaries aim to reconstruct the original input from its observation. In MIA, adversaries attempt to infer whether a specific data patch is included in the training datasets by analyzing the model’s behavior. To defend against these and other attacks, a widely adopted strategy is to inject artificial privacy noise into the local models to obscure sensitive information, through e.g., the well-known Gaussian mechanism [3]. To quantify the information leakage of each client’s contribution, local mutual information privacy (LMIP) and local differential privacy (LDP) are commonly employed as effective privacy measures that capture the information disclosure of a mechanism from different perspectives [4, 5, 6]. While a number of studies [7, 8] confirm the effectiveness of Gaussian mechanisms with provable convergence guarantees, their practical applicability remains limited by the inherent utility-privacy trade-off. Strong privacy requires substantial noise, whose accumulation at the server can severely impair model accuracy and hinder convergence. Conversely, maintaining stable convergence and satisfactory training performance requires the injected noise to remain sufficiently small. This inherent tension creates a fundamental conflict between convergence and privacy in FL.
The aforementioned issue can be effectively addressed through secure aggregation [9, 10], in which privacy noises (secret keys) are designed with coordinated structures such that they can be canceled out during the server aggregation. This approach preserves privacy while maintaining convergence speed, making secure aggregation an appealing solution to privacy-sensitive FL applications, where both privacy and model utility are preserved. However, the effectiveness of secure aggregation highly relies on stable communication, as the random link disruptions will break the carefully designed coordination of secret keys required for accurate reconstruction of the global model. Recent advances in dropout-resilient secure aggregation primarily build upon Shamir’s Secret Sharing (SSS) [9, 11], and the use of maximum distance separable (MDS) random variables in secret key construction [12]. While these techniques offer perfect privacy by guaranteeing zero information leakage, they encounter several limitations in practical FL scenarios. These include high encoding complexity and the lack of a clear, non-trivial mapping between real-valued model parameters and finite-field arithmetic, raising concerns about their practicality in real-world deployments.
@@ I-A Literature Review
Another challenge posed by unreliable communication is the persistent, strict sub-optimality of federated optimization. As the data generated at the network edge is inherently uncontrollable and typically follows non-independent and identically distributed (non-IID) patterns across clients, local models tend to deviate significantly from the global objective, known as client drift[13]. As a result, the global model may converge in unstable or misaligned directions. Unreliable communication further amplifies this issue by causing statistically uneven local contributions, potentially leading the model to a stationary point far from the true optimum. This phenomenon, known as objective inconsistency [14], results in sub-optimality and undermines model reliability. The presence of privacy noise further intensifies the problem: The injected noise will enlarge the step size, and when combined with distorted optimization trajectories caused by unreliable communication, it can prevent the global model from approaching the true global optimum, and may even cause divergence.
Existing approaches to mitigate communication stragglers include heterogeneity-aware client sampling [15], adaptive resource allocation[16], heterogeneity-aware aggregation [17, 18], collaborative relaying [19, 20], etc. But they all rely on precise prior information to ensure optimality, thus increasing the implementation complexity in practical deployments. In contrast, [21] eliminates this need by appropriate accumulation mechanisms. Further improvements are made to handle unknown client participation statistics[22] and to adapt to non-stationary participation patterns[23]. However, these methods are prone to divergence, as the accumulated gradient after several consecutive transmission failures can grow excessively. Another line of work applies coding strategies for straggler mitigation, eliminating the need for prior knowledge. For example, [24] applies the MDS code to reduce the outage frequency in FL, but a consistent objective is not guaranteed in heterogeneous networks. In contrast, gradient coding (GC) has emerged as an effective approach that guarantees the optimality of FL without requiring prior knowledge of network conditions. Particularly, GC employs cyclic gradient codes to enable exact recovery of the global model despite arbitrary communication patterns [25]. Several variations of GC have been proposed to address various practical constraints and scenarios [26, 27, 28, 29]. However, these methods typically impose heavy computational burdens due to dataset replication across clients. To overcome this limitation, [30] proposed cooperative GC (CoGC), a gradient-sharing framework to compute the exact sum, as shown in Fig. 1. The CoGC greatly improves the communication-computation-efficiency and ensures convergence to the optimal solution over heterogeneous networks without requiring prior knowledge of network statistics.
Any two: $\bm{\Theta}_{1}+\bm{\Theta}_{2}+\bm{\Theta}_{3}$ $\mathcal{D}_{1}$$\mathcal{D}_{2}$$\mathcal{D}_{3}$$\frac{1}{2}\bm{\Theta}_{1}+\bm{\Theta}_{2}$$\bm{\Theta}_{2}-\bm{\Theta}_{3}$$\frac{1}{2}\bm{\Theta}_{1}+\bm{\Theta}_{3}$$\bm{\Theta}_{1}$$\bm{\Theta}_{2}$$\bm{\Theta}_{3}$
Figure 1: An illustration of the CoGC framework [31] with $K=3$ clients, where all communication links are unreliable. Each client $k$ computes its local model $\bm{\Theta}_{k}$ based on its local datasets $\mathcal{D}_{k}$, and shares $\bm{\Theta}_{k}$ with $s=1$ neighbor. Then, clients compute and transmit the partial sums to the server.
I-B Motivations
On one hand, CoGC still confronts unresolved privacy risks, especially during the gradient-sharing phase. On the other hand, the binary nature of the GC decoding mechanism provides a promising solution to the problem of unreliable communication in secure aggregation, as it allows either the exact recovery of the global model or nothing meaningful at all in the presence of random user dropouts. This property ensures that the structure of aggregation remains intact whenever the number of received updates reaches the threshold. Following this intuition, we arrive at the first fundamental research question explored in this paper:
[…]
    Is it possible to keep exact recovery of the global model while preserving privacy globally, despite unreliable communication? (C1)
Since the coding structure enforces the coefficients in the final sums, this challenge can be addressed by incorporating secure aggregation into the CoGC framework with the careful design of the additive secret keys. However, flexibility in secret key design introduces a new issue: inconsistent and uneven privacy protection among clients. This leads to the second key research question, concerning fairness in privacy preservation:
[…]
    How to design secret keys to ensure fair privacy preservation among clients in secure aggregation? (C2)
We seek to systematically address these questions by formally defining them and analyzing them from the perspective of secret key construction. The significance of our work is three-fold: (i) We ultimately resolve the model privacy issues in CoGC at all levels, while preserving its original advantages; (ii) Meanwhile, the proposed scheme offers an alternative secure aggregation approach that is resilient to random client dropout, without sophisticated encoding or mapping operations or multiple rounds of communication with the server; (iii) We formally formulate and address the secure aggregation problem for FL in the real field, making our solution practical and directly applicable. Notably, we introduce and formalize fairness as a real-field privacy metric, providing both a clear motivation and a principled foundation.
I-C Our Contributions
We apply rigorous theoretical analysis and extensive experimental validations to support our findings. The key contributions are detailed below.
(C1) To achieve both resilience and secure aggregation against unreliable communication, we propose a novel Secure CoGC scheme, termed SecCoGC. The SecCoGC guarantees optimal convergence performance in heterogeneous federated networks, enables exact reconstruction of the global model at the server, and provides arbitrarily strong user privacy globally. The construction conditions and implementation procedures are tailored to be directly compatible with the most existing FL frameworks. Furthermore, the proposed construction ensures privacy protection against any decoding mechanisms, providing robust security against various model attacks.
(C2) To address fairness concerns in privacy preservation, we propose a Fair and Secure CoGC (Fair-SecCoGC) scheme that ensures equitable privacy protection across all participating clients, achieves optimal privacy under a total power constraint, and retains the full benefits of the original SecCoGC. We establish rigorous theoretical foundations along with a general construction approach for fair secret keys. In addition, we present an extremely simplified yet effective construction method that requires almost no computation. To the best of our knowledge, this work is the first to identify user fairness as a privacy metric in FL.
(C3) To ensure a thorough evaluation of model privacy, we provide multi-view and in-depth theoretical privacy analysis for SecCoGC and Fair-SecCoGC, using both LMIP and LDP, across intermittent networks and throughout all stages of the training process. These results, in turn, ensure that the key constructions are both theoretically sound and practically deployable. Furthermore, our analysis of real-field secure aggregation, where privacy noises are correlated, offers valuable insights and a foundation for future research.
(C4) To ensure a tractable analysis of the proposed methods, we evaluate the system reliability of SecCoGC and Fair-SecCoGC under arbitrary network conditions. Building upon this, we analyze their convergence behavior with an arbitrary choice of local solver in non-convex settings. Our theoretical results yield a standard convergence rate of $\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)$ with probability $99.74\%$ under arbitrary privacy levels. Notably, the convergence analysis also applies to CoGC and GC-based distributed training, while accommodating arbitrary local solvers for broader applicability.
I-D Paper Organization
The remainder of this paper is organized as follows. Section II introduces the system model and necessary preliminaries. Section III presents the formulation of secure aggregation in the real field together with the proposed SecCoGC scheme. Section IV extends this result by incorporating fairness concerns into the SecCoGC design. Privacy guarantees at each stage in FL are evaluated in Section V and Section VI, using LMIP and LDP, respectively. In Section VII, we analyze the system reliability of the proposed methods, while Section VIII examines its convergence behavior. The performance of the proposed methods is demonstrated through numerical comparisons with various benchmarks on multiple datasets in Section IX. Finally, conclusions are drawn in Section X, and supporting proofs are provided in the appendices.
@@ II System Model and Preliminaries
II-A Machine Learning (ML) Model
Consider a distributed system consisting of $K$ clients, each holding local datasets $\mathcal{D}_{k}$ on it, coordinated by a central server. The goal is to learn a global model $\bm{\Theta}\in\mathbb{R}^{D}$ that minimizes the global objective function (GOF) defined in (1):
@@ II-A Machine Learning (ML) Model
through iterative training and communication. The learning weight $w_{k}$ measures the importance of clients, satisfying $\sum_{k=1}^{K}w_{k}=1$. W.L.O.G., this paper assumes $w_{k}=\frac{1}{K}$. The local objective function (LOF) that measures the accuracy of the model $\bm{\Theta}$ on local datasets $\mathcal{D}_{k}$ is defined by the average loss on all training samples in $\mathcal{D}_{k}$, that is,
@@
II-B Communication Model
The inter-user and uplink communication channels are assumed to be orthogonal and modeled as independent Bernoulli r.v.s. Specifically, the link from client $k$ to the sever, denoted by $\tau_{k}$, is modeled by $\tau_{k}\sim\mathrm{Ber}(1-p_{k})$, the link from client $m$ to client $k$, denoted by $\tau_{k,m}$ is modeled by $\tau_{k,m}\sim\mathrm{Ber}(1-p_{k,m})$, where $p_{k}$ and $p_{k,m}$ are the probabilities when the channel is in outage. The $\tau_{k}=1$ and $\tau_{k,m}=1$ implies a successful communication and perfect recovery, instead, $\tau_{k}=0$ and $\tau_{k,m}=0$ implies a complete failure. For all $k$, $p_{k,k}=0$, $\tau_{k,k}=1$, since client $k$ has access to its own local model and requires no transmission. For $k\neq m$, if no transmission occurs from client $m$ to client $k$, then $p_{k,m}=1$, $\tau_{k,m}=0$, as clients do not have direct access to each other’s local models. Moreover, any link is statistically independent from each other, i.e., $\mathrm{(i)}\;\forall k_{1}\neq k_{2}:\tau_{k_{1}}\perp\tau_{k_{2}}$; $\mathrm{(ii)}\;\forall(k_{1},m_{1})\neq(k_{2},m_{2}):\tau_{k_{1},m_{1}}\perp\tau_{k_{2},m_{2}}$; and $\mathrm{(iii)}\;\forall k_{1},k_{2},m:\tau_{k_{1}}\perp\tau_{k_{2},m}$.
II-C Cyclic GC Codes
The GC codes consists of pair-wise cyclic matrices $\bm{G}$ and $\bm{C}$ that are carefully devised such that $\bm{C}\bm{G}=\mathbf{1}_{f\times K}$, where $f={K\choose s}$. In particular, $\bm{G}$ is the allocation matrix of size $K\times K$ in a cyclic pattern with $s$ nonzero entries per row, and $\bm{C}$ is the combination matrix of size $f\times K$, where each row represents a unique pattern with exactly $s$ zeros per row, encompassing all such possible patterns. The structure ensures that, even if up to $s$ rows are missing from $\bm{G}$, its resulting product with $\bm{C}$ still includes at least one all-one row. The construction for GC codes can be found in [25, Algo. II].
II-D CoGC Framework
Instead of replicating datasets across clients according to the nonzero patterns in $\bm{G}$ [25, Algo. II], CoGC enables gradient sharing among clients. Let $g_{k,m}$ denote the $(k,m)$-th element in $\bm{G}$, the following phases are carried out:

### v4

# Coding-Enforced Robust Secure Aggregation for Federated Learning Under Unreliable Communication

@@ Abstract
This work studies privacy-preserving federated learning (ppFL) under unreliable communication. In ppFL, zero-sum privacy noises enables privacy protection without sacrificing model accuracy, effectively overcoming the privacy-utility trade-off. However, in practice, unreliable communication can randomly disrupt the coordination of zero-sum noises, leading to aggregation errors and unpredictable partial participation, which severely harm the model accuracy and learning performance. To overcome these challenges, we propose a robust coding-enforced structured secure aggregation method, termed secure cooperative gradient coding (SecCoGC), which enables exact reconstruction of the global model under unreliable communication while allowing for arbitrarily strong privacy preservation. In this paper, a complete problem formulation and constructions of real-field zero-sum privacy noise are presented, and fairness is introduced as a privacy metric. Privacy across all protocol layers in SecCoGC is evaluated, accounting for the correlation among privacy noises and their linear combination under unreliable communication. Moreover, a distinct convergence analysis for the FL algorithm with a binary outcome for global model recovery is provided. Experimental results demonstrate that SecCoGC achieves strong resilience to unreliable communication while maintaining varying levels of privacy preservation, yielding test accuracy improvements of up to $20\%$-$70\%$ over existing benchmark methods.
@@ I Introduction
Federated learning (FL) enables numerous clients to collaboratively train models based on their local datasets, enhancing user privacy and reducing the volume of the transmitted data [1, 2]. However, the adversary can still infer sensitive information that is captured by local model updates, making FL susceptible to various model-based attacks, e.g., inversion attacks (IA) and membership inference attacks (MIA), etc.. To defend against these and other attacks, a widely adopted strategy in privacy-preserving FL (ppFL) is to inject privacy noise into local models to obscure sensitive information, through, e.g., the well-known Gaussian mechanism [3]. A number of studies confirm the effectiveness of independent Gaussian privacy noises with provable convergence guarantees [4, 5]. However, a strong privacy guarantee requires substantial noise at each client, leading to significant noise accumulation at the server, which severely impairs model accuracy and convergence. This inherent tension is called the utility-privacy trade-off. Both privacy leakage and model accuracy are significant concerns in FL. To mitigate the conflict between these two metrics, some researchers investigate the use of correlated privacy noises instead [6, 7]. Such designs allow for partial cancellation of randomness and mitigate the amplification issue of the accumulated privacy noise at the server, yet they are still constrained by the privacy-utility trade-off in nature.
Secure aggregation, which enables the server to learn the accurate aggregated model without observing the local ones, provides a promising solution to overcome the privacy-utility trade-off. Most studies on secure aggregation are based on homomorphic encryption, Shamir’s Secret Sharing (SSS), multi-party computation, zero-sum discrete and analog noises, and hybrid combinations of these techniques [8, 9, 10, 11, 12, 13, 14, 15, 16]  [^1] Zero-sum noise is considered a secure aggregation method, as it permits the server to obtain an accurate global model without revealing the local models [13, 14, 16]. . In these methods, the correct reconstruction of the global model requires that the coordinated encryption structure is reliable, which is highly dependent on reliable and predictable communication. However, in practice, communication channels are often unreliable due to real-world imperfections, e.g., fading and shadowing. Such communication uncertainty will result in random and unpredictable variations in both the pattern of participating clients and their cardinality, making the problem of secure aggregation particularly challenging under unreliable communication. Recent advances in dropout-resilient secure aggregation primarily build upon a combination of SSS with gradient coding (GC) on coded datasets [17], pairwise masking [9], and multi-round strategic secure communication scheme [13]. Despite their significant contributions, [17] incurs high encoding and decoding complexity, [9] allows the server to learn partial sums of the local models beyond the intended global model, [9, 13] suffers from the negative impact of partial participation, as will be discussed shortly.
@@
Another critical challenge in FL posed by unreliable communication is the persistent, strict sub-optimality arising from random partial participation. The data generated at the network edge is inherently uncontrollable and typically follows non-independent and identically distributed (non-IID) patterns across clients. For this reason, the local models tend to deviate largely from the global model, referred to as client drift[18]. Unreliable communication further intensifies this issue by introducing statistically uneven local contributions in the global aggregation, which guides the global model towards a stationary point far from the true optimum, referred to as objective inconsistency [19, 20].
In ppFL under unreliable communication conditions, the aforementioned phenomena become even more pronounced due to the mutual coupling between communication and privacy noise. First, the injected privacy noise will randomly perturb both the convergence direction and the effective step size, leading to an unstable optimization trajectory. When such instability interacts with unreliable communication, the resulting optimization trajectory becomes highly unpredictable and heterogeneous. Second, under unreliable communication, the variance of the aggregated noise becomes highly random and unstable. With the correlated privacy noise design, the aggregated noise at the server can grow substantially, as the correlated components fail to cancel effectively as intended. The combined effect of these facts can severely degrade the global model accuracy and impair training performance, and in extreme cases, may even cause divergence of FL.
Existing methods to mitigate the adverse effects of unreliable communication include heterogeneity-aware client sampling [21], adaptive resource allocation [20], heterogeneity-aware aggregation [22, 23], collaborative relaying [24, 25], etc. But they all rely on precise prior information to ensure their effectiveness. In contrast, [26, 27, 28] eliminates this need by appropriate accumulation mechanisms. However, they are prone to divergence, as the accumulated gradient norm can grow excessively after several consecutive transmission failures. Another line of work employs coding strategies to mitigate stragglers in FL without requiring prior information, including the usage of coded networks to reduce straggler occurrence [29], the combination of dataset-replication-based GC and coded datasets [17], and gradient-sharing-based cooperative gradient coding (CoGC) [30]. While extensive efforts have focused on enhancing robustness against unreliable communication, these methods generally leave privacy preservation unaddressed. As a result, a substantial gap remains between straggler-resilient design and robust privacy protection, highlighting a critical challenge that must be addressed.
I-A Our Contributions
To preserve both global model utility and user privacy under unreliable communication, this paper proposes a coding-enforced structured aggregation method, which simultaneously ensures the accurate computation of the global model and the complete cancellation of the zero-sum privacy noises, thereby significantly improving learning performance while allowing for arbitrarily strong privacy. Apart from this, the binary outcome inherent in the GC decoding mechanism provides an additional strong privacy guarantee, ensuring that the server aggregation is either intact or not meaningful. Both rigorous theoretical analysis and extensive experimental validations are applied to support our findings. The key contributions are summarized below.
(C1) A complete problem formulation  [^2] Notably, [16] presents an essential covariance-based construction for zero-sum noise in real field, which is advantageous in preset covariance and noise power. But a complete problem formulation is still lacking. of real-field zero-sum noise is present. To the best of our knowledge, this work first discusses the design principle, conditions, and constructions of real-field zero-sum noises. Moreover, it first introduces fairness  [^3] In this work, fairness is defined as allocation of privacy preservation levels in proportion to the clients’ learning weights [^4] The fairness concern can be further explored through algorithmic and system-level approaches, such as fairness-aware privacy budget allocation, adaptive client scheduling, and resource management strategies in FL system designs. These topics are left open for future research. as a performance metric in user privacy preservation. .
(C2) We propose a novel robust coding-enforced secure aggregation method, termed SecCoGC, which leverages coding techniques to enforce a determined aggregation structure over local models. SecCoGC remains resilient to unreliable communication in the following aspects:
  • The privacy noise is enforced to fully cancel out through the predetermined coded aggregation structure, regardless of its privacy level. This ensures the accuracy of the global model under unreliable communication, while allowing for arbitrarily strong privacy.
  • The global model is guaranteed to be formed correctly based on all local models, ensuring the optimality of FL algorithms and avoiding the negative effect arising from the unpredictable partial participation patterns of clients.
To the best of our knowledge, no prior work has applied coding schemes to ensure robust, secure aggregation in FL under unreliable communication  [^5] In [31], the authors introduce a secure computation scheme for linearly separable functions. While it shares some conceptual similarity with our approach, the focus and distributed learning setting differ significantly..
(C3)Multi-view and in-depth theoretical privacy analyses are provided for SecCoGC across all protocol layers under local differential privacy (LDP), covering both successful and failed global model reconstruction. Our analysis tackles four distinctive challenges:
  • In cooperative networks, a local update can be delivered to the server both directly and indirectly through neighboring nodes. This multi-path delivery introduces a substantial conceptual challenge in analyzing local differential privacy.
  • The analysis captures the correlation structure among privacy noises and their linear aggregation over intermittent networks, which poses significant challenges to LDP analysis, as the variance of the aggregated correlated noise becomes a random variable determined by network intermittency.
  • The correlation among secret keys enables inference based on a client’s secret key, and this leakage is properly accounted for in our analysis.
  • Unreliable communication leads to the simultaneous absence of the local model updates and the privacy noise, causing the sensitivity function to be unbounded.
To the best of our knowledge, no prior work has ever considered or tackled the above challenges. Our analysis offers valuable guidance for future privacy studies involving unreliable communication or cooperative networks.
(C4) A unique non-convex convergence analysis with an arbitrary choice of local solver is provided for SecCoGC and FL algorithms that result in binary recovery of the global model. Distinctively, both the number of consecutive training rounds between two adjacent successful reconstructions of the global model, and the total number of training rounds, are dependent random variables, which significantly complicates the convergence analysis. Our theoretical results yield a standard convergence rate of $\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)$ with probability $99.74\%$, holding under arbitrary local solvers and arbitrary privacy levels.
I-B Paper Organization
The remainder of this paper is organized as follows. Section II presents the system model and preliminaries. Section IV presents a complete formulation of real-field zero-sum noise. Section III introduces the proposed SecCoGC. Section V evaluates LDP across all layers in SecCoGC. Section VI analyses the convergence behavior of FL algorithms with binary global construction outcomes. Section VII provides extensive performance comparisons of the proposed method over multiple benchmarks. Conclusions are drawn in Section VIII, and supporting proofs are provided in the appendices of the supplementary file [32] [^6] The supplementary file can be found at https://arxiv.org/abs/2507.07565..
@@ II System Model and Preliminaries
II-A Federated Learning (FL)
Consider a FL system consisting of $K$ clients, each holding local datasets $\mathcal{D}_{k}$ on it, coordinated by a central server. The goal is to learn a global model $\bm{\Theta}\in\mathbb{R}^{D}$ that minimizes the global objective function (GOF) defined in (1):
@@ II-A Federated Learning (FL)
through iterative training and communication. The learning weight $w_{k}$ measures the importance of clients, satisfying $\sum_{k=1}^{K}w_{k}=1$. W.L.O.G., this paper assumes $w_{k}=\frac{1}{K}$. The local objective function (LOF) that measures the accuracy of the model $\bm{\Theta}$ on local datasets $\mathcal{D}_{k}$ is defined by the average loss on all training samples in $\mathcal{D}_{k}$,
@@
II-B Threat Model
Our threat model follows a standard assumption adopted in multi-party computation (MPC) [9, 33, 4]. The server and clients are assumed honest-but-curious: they faithfully execute the protocol yet may attempt to infer other local models based on publicly observable information and their local knowledge. To be more specific, the clients may try to learn the local model or the participation of a client by exploiting observable data, e.g., the received messages from its neighbors or the computed partial sums, and correlations among secret keys. The server may try to infer the local model, the participation of a client at another relaying client, or the participation of a client based on the combined partial sum using any predefined combinator.
II-C Communication Model
The inter-user and uplink communication channels are assumed to be orthogonal and modeled as independent Bernoulli r.v.s. The link $\tau_{k}$, from client $k$ to the sever, is modeled by $\tau_{k}\sim\mathrm{Ber}(1-p_{k})$, the link $\tau_{k,m}$, from client $m$ to client $k$, is modeled by $\tau_{k,m}\sim\mathrm{Ber}(1-p_{k,m})$, where $p_{k}$ and $p_{k,m}$ are the probabilities when the channel is in outage. The $\tau_{k}=1$ and $\tau_{k,m}=1$ implies a successful communication and perfect recovery, $\tau_{k}=0$ and $\tau_{k,m}=0$ implies a complete failure. For $\forall k$, $p_{k,k}=0$, $\tau_{k,k}=1$, since client $k$ has access to its own local model and requires no transmission. For $k\neq m$, if no transmission occurs from client $m$ to client $k$, then $p_{k,m}=1$, $\tau_{k,m}=0$, as clients do not have direct access to each other’s local models. Moreover, any link is assumed to be statistically independent from each other, i.e., $\mathrm{(i)}\;\forall k_{1}\neq k_{2}:\tau_{k_{1}}\perp\tau_{k_{2}}$; $\mathrm{(ii)}\;\forall(k_{1},m_{1})\neq(k_{2},m_{2}):\tau_{k_{1},m_{1}}\perp\tau_{k_{2},m_{2}}$; and $\mathrm{(iii)}\;\forall k_{1},k_{2},m:\tau_{k_{1}}\perp\tau_{k_{2},m}$.
II-D Cyclic Gradient Codes
The cyclic gradient code is formed by pair-wise cyclic matrices $\bm{G}$ and $\bm{C}$ that are devised such that $\bm{C}\bm{G}=\mathbf{1}_{f\times K}$, where $f={K\choose s}$. In particular, $\bm{G}$ is the allocation matrix of size $K\times K$ in a cyclic pattern with $s$ nonzero entries per row, and $\bm{C}$ is the combination matrix of size $f\times K$, where each row represents a unique pattern with exactly $s$ zeros per row, encompassing all such possible patterns. The structure ensures that, even if up to $s$ rows are missing from $\bm{G}$, its resulting product with $\bm{C}$ still includes at least one all-one row. The construction for GC codes can be found in [34, Algo. II].